how do i find the value of x is the following problem? 9/2 = x/4

Answer by Gator Tutoring(5)

Solve for x, 9/2 = x/4
9/2 = x/4
multiply both sides by 4
4 * 9/2 = 4 * x/4
36/2 = x
18 = x

Gator Tutoring
pichonz@gmail.com
Getting Math help couldn't get any easier.
How?
1. Email me all your questions (minimum of 5)
2. I post worked out solutions on Google/Amazon/Ebay
3. You purchase, I send file through email.
Costs:
Standard = Worked out solutions
Premium = Worked out solutions + strategies + definitions + comments. Answers from above are done using the premium service.
Standard > $1 per problem
Premium > $2 per problem
note: Problems with multiple questions are counted as separate problems.